OpenStudy (anonymous):

Francesca reads 57 pages every 2 h. How many pages can she read in 5 h?

OpenStudy (stefrheart):

First make a proportion

OpenStudy (anonymous):

ok

OpenStudy (stefrheart):

So what would it look like?

OpenStudy (anonymous):

i know you do cross multi and division but not sure howto really set it up,

OpenStudy (stefrheart):

well first lets put pages over hours

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Walkerman15

OpenStudy (stefrheart):

So it would look like 57/2 =x/3

OpenStudy (anonymous):

so im going to take 57x3 and then divide by 2?

OpenStudy (stefrheart):

yes 57*3=2x

OpenStudy (anonymous):

I came up with 85.50, does that sound right?

OpenStudy (stefrheart):

hmmm it could be 38..

OpenStudy (stefrheart):

Hold on

OpenStudy (stefrheart):

We could have either done it as a proportion or find how many pages per hour and then multiply it by hours

OpenStudy (anonymous):

Caleb deposited money into a savings account each week. The table shows the balance in his account after different numbers of weeks. Caleb graphed this information. What is the slope of the line that represents this situation? Number of weeks 3 6 9 15 Account balance ($) 135 270 405 675

OpenStudy (anonymous):

here is the other question @Walkerman15

OpenStudy (stefrheart):

You need to put them in fractions